OpenStudy (anonymous):

need help on a math work sheet these question are difficult please show work

OpenStudy (anonymous):

where are the problems/

OpenStudy (anonymous):

ah! marxin is back !

OpenStudy (anonymous):

did u expect me to stay away from this awesome site????

OpenStudy (anonymous):

nice worksheet!

OpenStudy (anonymous):

what is this from 1962?

OpenStudy (anonymous):

any specific problem?

OpenStudy (anonymous):

all of them startin with one i got to get these right

OpenStudy (anonymous):

Your last problem is wrong. It should be d. You foil the whole thing, then you add like terms. First Outer Inner Last

OpenStudy (anonymous):

and srry if i take a while cpu movin slow today

OpenStudy (anonymous):

first one is right second one is right third one is wrong. if you multiply out your selection you will see that you do not get 11x as the 'middle term"

OpenStudy (anonymous):

you would get 2x - 6x = -4x but you want 11x. i think a) will give it to you

OpenStudy (anonymous):

ok that seems right

OpenStudy (anonymous):

fourth one is wrong. \[(2t^2-t-28)\div (t-4)\] you are trying to solve \[2t^2-t-25=(t-4)(something)\]

OpenStudy (anonymous):

As soon as someone sees "difficult" on this website everyone flocks xDDD Whats up marxin?

OpenStudy (anonymous):

nuthin braw jus back with hm last day too get it done n yall r a big help

OpenStudy (anonymous):

that something has to look like \[(2t \pm b)\] because otherwise you will not get \[2t^2\] as your first term

OpenStudy (anonymous):

pick \[(2t+7)\] because \[2t\times t = 2t^2\] and \[-4\times 7 = -28\]

OpenStudy (anonymous):

I feel that. I hope it turns out good for you xP

OpenStudy (anonymous):

and unfortunately the fifth one is wrong because the expansion fo \[(a+b)^4=a^4+4a^3b+6a^2b^2+4ab^3+b^4\]

OpenStudy (anonymous):

so your coefficients should look like 1 4 6 4 1, not 1 5 10 5 1

OpenStudy (anonymous):

so its d right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

coefficients should be 1 4 6 4 1

OpenStudy (anonymous):

so then it's between a n b i go with a

OpenStudy (anonymous):

Its A. And it can't be the one with 1 5 10 5 1. Because those aren't correct. They would read: 1 5 10 10 5 1 :P And since it is (a-b)^4. You have alternating minus signs. So A.

OpenStudy (anonymous):

ok i get it we have to use coherants ok rady for the next work sheet